OpenStudy (anonymous):

A sum of money amounting to $4.40 consists of dimes and quarters. If there are 23 coins in all, how many are quarters?

OpenStudy (lgbasallote):

let x = dime y= quarter x + y = 23 10(cent)x + 25(cent)y = 440 cents x= 23 - y 10x + 25 y = 440 10(23 - y) + 25y = 440 230 -10y + 25y = 440 15y = 210 y = 14 quarters

OpenStudy (anonymous):

let a be dimes and b be quarters 0.1a+0.25b=4.40 a+b=23 a=23-b 2.3-0.1b+0.25b=4.40 0.15b=2.1 b=14 a=23-14 =9 There are 14 quarters and 9 dimes :)

OpenStudy (anonymous):

{d*10 + q*25 = 440, d + q = 23} d = 9, q = 14